### v1

# The Riemann Hypothesis in Oaxaca

The Riemann Hypothesis in Oaxaca
Carlos Segovia SECIHTI-UNAM-Oaxaca, México. e-mail: csegovia@im.unam.mx
Riemann [7] introduced the zeta function $\zeta(s)$ with the assertion that all of the nontrivial zeros have real part equal to $1/2$. Later, Robin [8] established an equivalence between the RH and the inequality:
$\frac{\sigma(n)}{n}<e^{\gamma}\log\log n,\quad\textrm{ for }n>5040\,$, (1)
with $\sigma(n)$ the sum of divisors and $\gamma$ the Euler–Mascheroni constant. Lagarias [4] provided an elementary equivalence of the RH with the inequality:
$\sigma(n)\leq H_{n}+e^{H_{n}}\log H_{n}\,$, (2)
where $H_{n}$ is the $n$-Harmonic number. In the spring of 2025, the student J. Damian Espinosa [3] presented to me a simple equivalence of the RH with the inequality:
$A(n):=\sum_{k=0}^{\infty}a_{k}\frac{(H_{n})^{k}}{k!}\geq\sigma(n)\,$, (3)
for $H_{k}-\gamma\leq a_{k}\leq\log(k)+\frac{1}{k}$, $k\geq 1$.
Proof.
Take the sum
$\sum_{k=0}^{\infty}\left(\frac{H_{k}-\gamma}{k!}\right)x^{k}$ $=\sum_{k=1}^{\infty}H_{k}\frac{x^{k}}{k!}-\gamma e^{x}\,$, (4)
$=e^{x}(\log x+\gamma+E_{1}(x))-\gamma e^{x}\,$, (5)
$=e^{x}(\log x+E_{1}(x))\,$, (6)
where we use
$\sum_{k=1}^{\infty}H_{k}\frac{x^{k}}{k!}$ $=\int_{0}^{1}\frac{e^{x}-e^{xt}}{1-t}dt\,$, (7)
$=e^{x}\int_{0}^{x}\frac{1-e^{-t}}{t}dt\,$, (8)
which is equal to $e^{x}(\log x+\gamma+E_{1}(x))$ by [1, 5.1.39]. By Lagarias [4], for $n\geq 3$,
$e^{\gamma}n\log\log n\leq e^{H_{n}}\log H_{n}\,.$ (9)
Suppose the RH, by Robin [8], for $n>5040$, we have
$\sigma(n)\leq e^{\gamma}n\log\log n\,$, (10)
For $1\leq n\leq 5040$ we can verify by computer that $\sigma(n)\leq A(n)$.
Suppose the inequality $\sigma(n)\leq A(n)$. If the RH is false, by Robin [8], there exists constants $0<\beta<\frac{1}{2}$ and $C>0$ such that the inequality
$\sigma(n)\geq e^{\gamma}n\log\log n+\frac{Cn\log\log n}{(\log n)^{\beta}}\,$, (11)
holds for infinitely many $n$. In addition, by Lagarias [4], for $n\geq 3$, we have the inequality
$e^{H_{n}}\log H_{n}\leq e^{\gamma}n\log\log n+\frac{3e^{\gamma}n}{2\log n}\,.$ (12)
Because of (6), we obtain
$\frac{Cn\log\log n}{(\log n)^{\beta}}$ $\leq\frac{3e^{\gamma}n}{2\log(n)}+e^{H_{n}}E_{1}(H_{n})\,$, (13)
$\leq\frac{3e^{\gamma}n}{2\log n}+\frac{e^{\gamma}(n+1)}{\log n}\,$, (14)
$=\frac{e^{\gamma}}{\log n}\left(\frac{5}{2}n+1\right)\,$, (15)
where $e^{H_{n}}E_{1}(H_{n})\leq e^{H_{n}}/H_{n}\leq e^{\gamma}(n+1)/\log n$ since $\log n<H_{n}<\log(n+1)+\gamma$ for $n\geq 1$.
Therefore, we have the inequality:
$\frac{C}{e^{\gamma}}\log\log n\leq(\log n)^{\beta-1}\left(\frac{5}{2}+\frac{1}{n}\right)\,.$ (16)
Taking limits with $n\rightarrow\infty$, we obtain the left-side tend to infinity, while the right-side tend to zero, which is a contradiction. We can increase the coefficients $H_{k}-\gamma\leq a_{k}\leq\log k+1/k$ providing the following inequality (see de Appendix):
$A(n)\leq e^{H_{n}}\log H_{n}+\frac{2e^{\gamma}(n+1)}{\log n}\,.$ (17)
This inequality implies we do not exceed the threshold $\lim_{n\rightarrow\infty}A(n)/n\log\log n=e^{\gamma}$. ∎
The Young lattice is one of the most important and beautiful structures in mathematics. The connection with the RH arises through the interpretation of $A(n)$ in terms of integer partitions:
$\exp(x_{1}+\cdots+x_{n})=\sum_{j=1}^{\infty}\sum_{\mu\vdash j}\frac{m_{\mu}\left(x_{1},x_{2}\ldots,x_{n}\right)}{\mu_{1}!\,\mu_{2}!\,\ldots\,\mu_{\ell(\mu)}!}\,$, (18)
for $m_{\mu}$, the monomial symmetric polynomial associated with the partition $\mu=[\mu_{1},\cdots,\mu_{l(\mu)}]$. Consider partitions of $j$ with order $i$ and define the sums:
$E_{i,j}(n):=\sum_{\mu\vdash j,\,\ell(\mu)=i}\frac{m_{\mu}(x_{1},\dots,x_{n})}{\mu_{1}!\mu_{2}!\dots\mu_{i}!}\,.$ (19)
We denote the $r$-sum as follows:
$A_{r}(n)=\sum_{i=1}^{n}\log(i+r)E_{i,i+r-1}(n)\,$, (20)
and we obtain $A(n)=\sum_{r=1}^{\infty}A_{r}(n)$ with $x_{i}=1/i$, for $i=1,\ldots,n$.
The case $r=1$ is a well understood logarithmic structure
$A_{1}(n)=\sum_{m=1}^{n}\log(m+1)e_{m}(n)\,$, (21)
where the $e_{m}$ are the elementary symmetric polynomials with
[…]
the numerator is the sign-less Stirling number of first kind, see [2]. Denote by $C_{n}$ the number of cycles of a permutation of $n$ elements. The elementary symmetric polynomials have the following probabilistic interpretation:
$e_{m}(n)=(n+1)\mathbb{P}(C_{m+1}=m+1)\,$, (23)
where $\mathbb{P}(C_{n+1}=m+1)$ denotes the probability that a permutation of $n+1$ elements, chosen uniformly at random, has exactly $m+1$ cycles; see [2].
Consequently,
$A_{1}(n)=(n+1)\sum_{k=1}^{n+1}\log k\cdot\mathbb{P}(C_{n+1}=k)=(n+1)\cdot\mathbb{E}[\log C_{n+1}]\,.$ (24)
For $r\geq 2$, we found the following determinantal formulas:
$p_{k}=\begin{vmatrix}e_{1}&1&0&0&\cdots&0\\ 2e_{2}&e_{1}&1&0&\cdots&0\\ 3e_{3}&e_{2}&e_{1}&1&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ ke_{k}&e_{k-1}&e_{k-2}&\cdots&e_{2}&e_{1}\end{vmatrix}\,,\hskip 28.45274ptm_{[r,1^{l}]}=\begin{vmatrix}e_{1}&1&0&0&\cdots&0\\ 2e_{2}&e_{1}&1&0&\cdots&0\\ 3e_{3}&e_{2}&e_{1}&1&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ (r-1)e_{r-1}&e_{r-2}&e_{r-3}&\cdots&e_{1}&1\\ (l+r)e_{l+r}&e_{l+r-1}&e_{l+r-2}&\cdots&e_{l+2}&e_{l+1}\end{vmatrix}\,$, (25)
[…]
We denote the sub-summand of $A_{r}(n)$:
$\tilde{A}_{r}(n)=\frac{1}{r!}\sum_{l=0}^{n-1}\log(l+r+1)\,m_{[r,1^{l}]}(x_{1},x_{2},\ldots,x_{n})\,.$ (26)
The use of the previous determinantal formulas give the expresion:
$\tilde{A}_{r}(n)=\frac{1}{r!}\det\begin{pmatrix}e_{1}(n)&1&0&\cdots&0\\ 2e_{2}(n)&e_{1}(n)&1&\cdots&0\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ (r-1)e_{r-1}(n)&e_{r-2}(n)&\cdots&e_{1}&1\\ S_{1}^{(r)}&S_{2}^{(r)}&\cdots&S_{r-1}^{(r)}&S_{r}^{(r)}\end{pmatrix}\,$, (27)
where
$S^{(r)}_{1}=\sum_{m=1}^{n}\log(m+r)\cdot(m+r-1)e_{m+r-1}(n)$ (28)
and
$S^{(r)}_{j}=\sum_{m=1}^{n}\log(m+r)\cdot e_{m+r-j}(n),\quad j=2,\ldots,r\,.$ (29)
Remember $e_{m}(n)=(n+1)\mathbb{P}(C=m+1)$, $m\geq 0$. For $C:=C_{n+1}$, denote the following:
$E_{1}=\mathbb{E}[(C-1)\log C]\textrm{ and }E_{j}=\mathbb{E}[\log(C+j-1)],\textrm{ for }j\geq 2\,.$ (30)
Consequently, we write
$S_{1}^{(r)}=(n+1)E_{1}-(n+1)\sum_{t=1}^{r}\log t\cdot(t-1)\,\mathbb{P}(C=t)=(n+1)E_{1}-\sum_{t=1}^{r}\log t\cdot(t-1)\,e_{t-1}(n)$ (31)
and for $j\geq 2$,
$S_{j}^{(r)}=(n+1)E_{j}-(n+1)\sum_{t=1}^{r-j+1}\log(t+j-1)\,\mathbb{P}(C=t)=(n+1)E_{j}-\sum_{t=1}^{r-j+1}\log(t+j-1)\,e_{t-1}(n),\quad j\geq 2\,.$ (32)
where $e_{t-1}(n)=(n+1)\mathbb{P}(C=t)$ states our previous notation.
Set $\Delta_{1}=-\sum_{t=1}^{r}\log t\cdot(t-1)e_{t-1}$ and $\Delta_{j}=-\sum_{t=1}^{r-j+1}\log(t+j-1)e_{t-1}$, for $j\geq 2$. Because the matrix below has zero determinant:
$\begin{pmatrix}e_{1}&1&0&0&\cdots&0\\ 2e_{2}&e_{1}&1&0&\cdots&0\\ 3e_{3}&e_{2}&e_{1}&1&\cdots&0\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ (r-1)e_{r-1}&e_{r-2}&e_{r-3}&\cdots&e_{1}&1\\ \Delta_{1}&\Delta_{2}&\Delta_{3}&\cdots&\Delta_{r-1}&\Delta_{r}\end{pmatrix}\,$, (33)
therefore, we obtain the formula, for $r\geq 2$:
$\tilde{A}_{r}(n)=\frac{n+1}{r!}\sum_{j=1}^{r}(-1)^{r+j}p_{j-1}(n)E_{j}\,.$ (34)
In order to find expressions for $E_{j}$, $1\leq j\leq r$, we have the following:
$\mathbb{E}\bigl[\log(C_{n}+a)\bigr]=\log\bigl(H_{n}+a\bigr)+\sum_{m=2}^{\infty}\frac{(-1)^{m-1}}{m}\frac{\mathbb{E}[(C_{n}-H_{n})^{m}]}{(H_{n}+a)^{m}},$ (35)
and
$\mathbb{E}[C_{n}\log C_{n}]=H_{n}\log H_{n}-\sum_{m=2}^{\infty}\frac{(-1)^{m-1}}{m(m-1)}\frac{\mathbb{E}[(C_{n}-H_{n})^{m}]}{H_{n}^{m-1}}\,.$ (36)
The central moments $\mu_{m}:=\mathbb{E}[(C-H_{n})^{m}]$ are given in terms of the cumulants of the number of cycles of permutations $C_{n}$:
[…]
where $S(m,\ell)$ is the Stirling of second kind. The central moments $\mu_{m}=B_{m}(0,\kappa_{2},\ldots,\kappa_{m})$ are the complete Bell polynomials with positive integer coefficients evaluated in the cumulants:
$\mu_{1}=0,\quad\mu_{2}=\kappa_{2},\quad\mu_{3}=\kappa_{3},\quad\mu_{4}=\kappa_{4}+3\kappa_{2}^{2},\quad\mu_{5}=\kappa_{5}+10\kappa_{3}\kappa_{2},\quad\mu_{6}=\kappa_{6}+15\kappa_{4}\kappa_{2}+10\kappa_{3}^{2}+15\kappa_{2}^{3},\ldots$ (38)
It is not difficult to convince ourself that $B_{m}(0,\kappa_{2},\ldots,\kappa_{m})\sim c(\log n)^{\lfloor m/2\rfloor}$ for $c$ a constant. In addition, we use $H_{n+1}=\log n+\gamma+O(1/n)$ and $\log H_{n+1}=\log\log n+\gamma/\log n+o(1/\log n)$ and we obtain:
$E_{1}$ $=\mathbb{E}[C\log C]-\mathbb{E}[\log C]\,$, (39)
$=\left[H_{n+1}\log H_{n+1}+\frac{\kappa_{2}}{2H_{n+1}}+O\left(\frac{1}{\log n}\right)\right]-\left[\log H_{n+1}+O\left(\frac{1}{\log n}\right)\right]\,$, (40)
$=\log H_{n+1}(H_{n+1}-1)+\frac{\kappa_{2}}{2H_{n+1}}+O\left(\frac{1}{\log n}\right)\,$, (41)
$=\left(\log\log n+\frac{\gamma}{\log n}+o\left(\frac{1}{\log n}\right)\right)\left(\log n+\gamma-1+O\left(\frac{1}{n}\right)\right)+O(1)\,$, (42)
$=\log n\log\log n+(\gamma-1)\log\log n+\gamma+O(1)\,$, (43)
and
$E_{j}$ $=\mathbb{E}(\log(C+j-1))\,$, (44)
$=\left[\log H_{n+1}+\frac{j-1}{H_{n+1}}+o\left(\frac{1}{H_{n+1}^{2}}\right)\right]+O\left(\frac{1}{\log n}\right)\,$, (45)
$=\log\log n+O(1)\,$, (46)
for $j\geq 2$. As a consequence, we have the following:
$H_{n}E_{2}-E_{1}=\log\log n-\gamma+O(1)\,.$ (47)
We deduce the following limits:
$\rho_{1}=\lim_{n\rightarrow\infty}\frac{A_{1}(n)}{n\log\log n}=1,\quad\rho_{2}=\lim_{n\rightarrow\infty}\frac{A_{2}(n)}{n\log\log n}=\lim_{n\rightarrow\infty}\frac{(n+1)}{2n}\frac{H_{n}E-E_{1}}{\log\log n}=\frac{1}{2}\,$, (48)
and for $r\geq 2$, we have:
$\rho_{r}:=\lim_{n\rightarrow\infty}\frac{\tilde{A}_{r}(n)}{n\log\log n}=\lim_{n\rightarrow\infty}\frac{(n+1)(-1^{r})}{r!n}\left(\frac{-E_{1}+p_{1}(n)E_{2}-p_{2}(n)E_{3}+\ldots+(-1)^{r}p_{r-1}(n)E_{r}}{\log\log n}\right)\,.$ (49)
Therefore, we obtain
$\rho_{r}=\frac{1}{r!}\left[(-1)^{r}+\sum_{j=3}^{r}(-1)^{r+j}\zeta(j-1)\right]\,.$ (50)
From the proof of the RH equivalence we know that $\lim_{n\rightarrow\infty}A(n)/n\log\log n=e^{\gamma}\approx 1.7810\ldots$ The previous formula provides the proportions:
[…]
Nevertheless, we get a strict inequality $1.6359<1.7810$ where we need to add the sub-summands for partitions which are not of the form $r,1^{l}]$.
For instance, for $r=3$, we have those partitions associated to $[2,2]$ and for $r=4$, we have to add the partitions associated to $[3,2]$ and $[2,2,2]$. The correction term for the partitions associated to $[2,2]$, has the following form:
$R_{3}(n)=\frac{1}{4}\sum_{l=0}^{n-2}\log(l+5)m_{[2,2,1^{l}]}=\frac{1}{4}\sum_{l=0}^{n-2}\log(l+5)\left[e_{l+2}e_{2}-(l+2)e_{l+3}e_{1}+\frac{(l+1)(l+4)}{2}e_{l+4}\right]\,.$ (51)
Computational calculations provides the proportion between $A_{r}(n)$ and $R_{r}(n)=A_{r}(n)-\tilde{A}_{r}(n)$ increases at the beginning but at some moment it should decrease, leaving the question of when this occurs and how.
Our interest comes from speculative arrangements of the divisors: positive integers $n$ with enough divisors, not a perfect square, we enumerate the divisors below the square root,
$1=d_{1}<d_{2}<\cdots<d_{k}<\sqrt{n}$ (52)
and the sum of divisors satisfies:
$\sigma(n)=\sum_{i=1}^{k}d_{i}+\frac{n}{d_{i}}$ (53)
For the first summand $A_{1}(n)\approx n\log\log n$, we consider all the sum of divisors in pairs with divisors $d_{i}\leq d^{(1)}(n)$, where $d^{(1)}(n)$ denotes the largest divisor satisfying the inequality:
$\sum_{d_{i}|n}^{d^{(1)}(n)}d_{i}+\frac{n}{d_{i}}\leq n\log\log n$ (54)
For colossally abundant numbers A004490, superabundant numbers A004394 and almost for the highly abundant numbers A002093, we have $d^{(1)}(n)\approx\lceil\gamma\log n\rceil$. At the moment we are not able to give a proof for colossally abundant numbers. However, a proof of the RH in this direction must have this inequality as first step with $d^{(1)}(n)\approx\lceil\gamma\log n\rceil$.
For the second summand $A_{2}(n)\approx\frac{1}{2}n\log\log n$, we do not have a explicit approximation of $d^{(2)}(n)$ in terms of a formula of logarithms with known constants.
$n$ $d^{(1)}(n)$ $\gamma\log n$ $d^{(2)}(n)$
2 0 0.400095410701532 0
6 1 1.03423163337409 1
12 1 1.43432704407562 2
60 2 2.36331981881901 4
120 3 2.76341522952054 5
360 3 3.39755145219309 8
2520 4 4.52076127271870 14
5040 5 4.92085668342023 16
55440 6 6.30495939767278 30
720720 8 7.78548834647870 44
1441440 8 8.18558375718024 48
4324320 9 8.81971997985279 56
21621600 9 9.74871275459618 72
367567200 11 11.3840878787935 91
6983776800 13 13.0836641819150 112
160626866400 14 14.8935205605575 140
321253732800 15 15.2936159712590 150
9316358251200 17 17.2372718726848 187
288807105787200 19 19.2194230801851 225
2021649740510400 20 20.3426329007107 248
6064949221531200 20 20.9767691233832 266
224403121196654400 22 23.0610475072350 310
9200527969062830400 25 25.2045794768775 368
395622702669701707200 27 27.3756031024853 418
791245405339403414400 27 27.7756985131868 430
37188534050951960476800 29 29.9980640210769 483
1970992304700453905270400 31 32.2897787078111 555
116288545977326780410953600 34 34.6433971946560 629
581442729886633902054768000 35 35.5723899693994 660
35468006523084668025340848000 37 37.9452507602346 740
Appendix
Take the coefficient $a_{k}=\log(k)+\frac{1}{k}$ we have the inequality:
$\sum_{k=1}^{\infty}\left(\log k+\frac{1}{k}\right)\frac{(H_{n})^{k}}{k!}\leq e^{H_{n}}\log H_{n}+\frac{2e^{\gamma}(n+1)}{\log n}\,$, (55)
for $n>1$.
Proof.
We have the following sequence of identities and inequalities:
$\sum_{k=1}^{\infty}\left(\log(k)+\frac{1}{k}\right)\frac{(H_{n})^{k}}{k!}$ $=\sum_{k=1}^{\infty}\log(k)\frac{(H_{n})^{k}}{k!}+\sum_{k=1}^{\infty}\frac{(H_{n})^{k}}{k\cdot k!}\,$, (56)
$<\sum_{k=1}^{\infty}(H_{k}-\gamma)\frac{(H_{n})^{k}}{k!}+E_{i}(H_{n})-\gamma-\log(H_{n})\,$, (57)
$=e^{H_{n}}\log(H_{n})+e^{H_{n}}E_{1}(H_{n})+E_{i}(H_{n})-\log(H_{n})\,.$ (58)
where we used [1, 5.1.10] together with $\log(k)<H_{k}-\gamma$, for $k\geq 1$, and (6). Consequently, it rest to show
$e^{H_{n}}E_{1}(H_{n})+E_{i}(H_{n})\leq\log(H_{n})+\frac{2e^{\gamma}(n+1)}{\log(n)}\,.$ (59)
We have the well-known inequality $e^{x}E_{1}(x)\leq 1/x$ and, we can show, for $x>2$, the inequality $E_{i}(x)\leq 2e^{x}/x$. In addition, $e^{H_{n}}/H_{n}\leq e^{\gamma}(n+1)/\log(n)$ since $\log(n)<H_{n}<\log(n+1)+\gamma$ for $n\geq 1$. In addition, $\log H_{n}>1/\log n$, for $n>3$. Thus (59) follows for $H_{n}>2$, the lower cases can be verified. ∎

### v2

@@
 The Riemann Hypothesis in Oaxaca
-The Riemann Hypothesis in Oaxaca
-Carlos Segovia SECIHTI-UNAM-Oaxaca, México. e-mail: csegovia@im.unam.mx
+Abstract
+An equivalence of the Riemann Hypothesis (RH) enables a direct bridge to the Young lattice.
+In specific, the classical threshold $\lim_{n\to\infty}\sigma(n)/(n\log\log n)=e^{\gamma}\approx 1.78107$, derived from the asymptotic behavior of the sum-of-divisors function, admits a combinatorial interpretation via limiting proportions associated to specific families of integer partitions.
 Riemann [7] introduced the zeta function $\zeta(s)$ with the assertion that all of the nontrivial zeros have real part equal to $1/2$.
@@
 $\sigma(n)\leq H_{n}+e^{H_{n}}\log H_{n}\,$, (2)
-where $H_{n}$ is the $n$-Harmonic number. In the spring of 2025, the student J. Damian Espinosa [3] presented to me a simple equivalence of the RH with the inequality:
+where $H_{n}$ is the $n$-Harmonic number. Recently, Espinosa published an equivalence of the RH in [3]. In what follows I present an sketch of his proof:
 $A(n):=\sum_{k=0}^{\infty}a_{k}\frac{(H_{n})^{k}}{k!}\geq\sigma(n)\,$, (3)
@@
 $E_{i,j}(n):=\sum_{\mu\vdash j,\,\ell(\mu)=i}\frac{m_{\mu}(x_{1},\dots,x_{n})}{\mu_{1}!\mu_{2}!\dots\mu_{i}!}\,.$ (19)
+Recently, Espinosa in [4], provides an interesting identity summing appropriately the monomials in $E_{i,j}(n)$ called the assembly theorem.
+The following is an improvement adding the coefficients in order to obtain complete sums.
 We denote the $r$-sum as follows: